# Equitable Allocation for Mixtures of Goods and Chores

### v2

@@ Appendix A Omitted Details of Section 3
(continued) We exhibit a reduction from $2$-partition, where given a multiset $U=\{b_{1},b_{2}\ldots b_{m}\}$ of integers with sum $2T$, the task is to decide if there is a partition into two subsets $S$ and $U\setminus S$ such that sum of the numbers in both the partitions equals $T$. We construct the reduced allocation instance as follows. We create $4$ agents, $m$ set-items $\{o_{1},\ldots,o_{m}\}$, and $4$ dummy items $\{d_{1},\ldots d_{4}\}$. The first two agents value the set-items at $\{b_{1},\ldots,b_{m}\}$ and all the dummy items at $-3T$. The last two agents value the set items at 00 and the dummy items at $T$. This completes the construction (See Table 2.) We now argue the equivalence.
Forward Direction.
Reverse Direction.
Suppose there is an EQ1 allocation, say $A$, under the reduced instance. Then, note that $A$ assigns at most one dummy item to the agents $a_{1}$ and $a_{2}$, each. If not, say WLOG, $a_{1}$ receives $\{d_{1},d_{2}\}$ under $A$, then the maximum utility $a_{1}$ can derive is $-4T$ (where it gets all the set items as well). But, $a_{1}$ violates EQ1 with respect to $a_{3}$ and $a_{4}$, whose minimum utility is 00 each, and $a_{1}$ derives negative utility even if it removes the item $d_{1}$ from its bundle. Therefore, $a_{1}$ and $a_{2}$ can get at most one dummy item each. We now consider the following cases:
  • $a_{1}$ receives $d_{1}$ and $a_{2}$ does not receive any dummy item. Then, WLOG, $a_{3}$ gets one dummy item and $a_{4}$ gets two dummy items. To be consistent with EQ1 against $a_{4}$, agent $a_{1}$ must get all the set items, thereby deriving a utility of $-3T+2T$. This leaves $a_{2}$ empty-handed and hence, it violates EQ1 with respect to $a_{4}$. Therefore, since $A$ is EQ1, this case does not arise.
This concludes the argument. ∎
See 3.3
Proof.
The algorithm iteratively picks a poor agent—breaking ties arbitrarily—and allocates it the most valuable good among the remaining items in $O^{+}$ according to its preference. Once $O^{+}$ is exhausted, it picks a rich agent who then receives her most disliked item from $O^{-}$.
@@
We show that Algorithm 1 returns an EQ1 allocation when the instance has $\{-1,1\}$ normalized valuations. First, note that since the instance is normalized, the reduced instance restricted to items in only $O^{\pm}$ is also normalized. That is, every agent assigns a value of $1$ to exactly $k_{1}$ items from $O^{\pm}$ and a value of $-1$ to exactly $k_{2}$ items from $O^{\pm}$, where $k_{1}$ and $k_{2}$ are constants. The idea is to first allocate all the items in $O^{\pm}$ in an almost equitable way and then to extend the partial EQ1 allocation to a complete allocation using Lemma 3.5. To that end, we first show that there exists a partial EQ1 allocation that allocates all items in $O^{\pm}$.
Lemma A.1.
@@
We first describe the algorithm. We iteratively pick a poor agent and allocate it an item it values at $1$ (a good) from $O^{\pm}$. If there are no such items in $O^{\pm}$, then we iteratively pick one of the rich agents and allocate it an item it values at $-1$ (a chore) from $O^{\pm}$. At this point, it is easy to see that the partial allocation is EQ1. We denote the set of agents with minimum utility as poor agents $(P)$ and those with maximum utility as rich agents ($R$). Now suppose that the remaining items in $O^{\pm}$ are all chores for the poor agents and goods for the rich agents, and if they are allocated to any of them, they increase the amount of inequity and hence may violate EQ1. Therefore, in order to move towards an EQ1 allocation that allocates all the items in $O^{\pm}$, we now aim to convert a poor (rich) agent to rich (poor) by re-allocating one of the allocated items, so that the converted agent can now be a potential owner of one of the remaining unallocated items while maintaining EQ1. We do so by executing one of the following transfers at a time, each of which allows us to allocate at least one unallocated item while respecting EQ1.
@@
    Every $p$ values every good in $A_{r}$ at $-1$.
@@
    Every $r$ values every chore in $A_{p}$ at $1$
@@
    Every $p$ values every chore in $A_{p^{\prime}}$ at $-1$.
With these valuations at hand, we are now ready to prove Lemma A.1.
Proof of Lemma A.1.
@@
    If $|P|>|R|$, then we take a good ($1$-valued item) from $A_{r}$ and allocate it to $p$ who necessarily values it at $-1$ (else, there is an R-P transfer). We do this for $|R|$ disjoint pairs of a rich and a poor agent, thereby decreasing the utility of all these pairs by $-1$. But since $|P|>|R|$, we have a poor agent $p^{\prime}$ whose value remains intact, which in turn, makes him one of the rich agents after the above transfers. Now, the remaining unallocated item, which was a chore for $p^{\prime}$ can be allocated to it without violating EQ1.
@@
    Even if $p$ values every good in every other poor agent’s bundle at $1$, then also only $|P|k$ out of $|R|(k+1)+1$ such items are accounted for. (Note that $p$ can not value a chore in $A_{p^{\prime}}$ at $1$, else there exists a P-P transfer). Also, note that $p$ cannot value any of the $|R|(k+1)$ items allocated to the rich agents at $1$, else there is an R-P transfer. This implies that there must be at least $(|R|-|P|)k+|R|+1$ many allocated items valued at $1$ by $p$, outside of the $|R|(k+1)$ and $|P|k$ items allocated as goods to the rich and poor agents respectively. At this point, we say that $r$ stays ahead of $p$ in terms of the number of items valued as a good.
    Since $v_{r}(A_{r})=k+1$, in particular, $r$ derives its value of $k+1$ from a set of $k+1$ items valued at $1$ by itself. We call these $k+1$ items as contributing goods in the bundle $A_{r}$. There can be additional items in $A_{r}$, however, they do not contribute to the utility of $r$. Any such additional good $g$ (chore $c$) in $A_{r}$ is non-contributing, that is, it can be paired up with a chore $c$ (good $g$) in $A_{r}$ so that the final utility contribution of $g\cup c$ to $r$ is 00. We refer to these pairs as non-contributing good-chore pair $(g,c)\in A_{r}$. Likewise, for any poor agent $p$, there are $k$ contributing goods in $A_{p}$, and the rest of the goods (chores) in $A_{p}$ can be paired up with a chore (good) in $A_{p}$ so their cumulative utility contribution is 00.
@@
    For $p$ to compensate for this deficit, it must value at least $(|R|-|P|)k+|R|+1$ many allocated items, outside of the non-contributing pairs in any rich bundle, at $1$. Let’s call this set of goods for $p$ as $E_{g}^{p}$. Since we have counted for the contributing goods in any poor bundle (at most $|P|k$ of them), for any $g\in E_{g}^{p}$, we must have that it is allocated to a poor agent $p^{\prime}$ as a non-contributing pair. (Again, the non-contributing pairs allocated to $r$ are already counted for). Let $c$ be the corresponding non-contributing chore in $A_{p^{\prime}}$. If $v_{p}(c)=1$, then we have a P-P transfer. So, for all $p\in P$, we have $v_{p}(c)=-1$. Now, if any rich agent values $c$ at $1$, then it again stays ahead of the poor agents by this additional $1$ valued item (which is outside its bundle), and the process of compensating for the deficit does not terminate. Otherwise, $v_{r}(c)=-1$, but this is a contradiction to the fact that $c\in O^{\pm}$.
@@
$r_{1}$ $1$ $1$ $1$ $1$ $1$ $1$ $\boxed{1,-1}$ $-1$,$-1$ $1$ $1$
$r_{2}$ $1$ $1$ $1$ $1$ $1$ $1$ $-1$,$-1$ $1$ $1$
$p_{1}$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$ $1$ $1$ $1$ $1$ $-1,\star$ $1$, $-1$ $-1$ $-1$
$p_{2}$ $-1$ $-1$ $-1$ $-1$ $-1$ $-1$ $1$ $1$ $1$ $1$ $-1,\star$ 1,$1$ , - 1 $-1$ $-1$
Table 5: A schematic of Case (2b) in the proof of Lemma A.1. $r_{1}$ and $r_{2}$ are two rich agents with utility $3$ each and $p_{1}$ and $p_{2}$ are two poor agents with utility $2$ each. The values in gray are forced, otherwise, there is a feasible transfer. The values in red depict the contradiction that there is an item $c$ such that $c\notin O^{\pm}$
@@
Lemma A.1 gives us a partial EQ1 allocation $A$. Now we can allocate the items from $O^{+}$ to a poor agent iteratively and once $O^{+}$ is exhausted, we allocate the item from $O^{-}$ to a rich agent iteratively (Lemma 3.5). This ensures that $A$ is EQ1 at every step of the algorithm. Consequently, the complete allocation $A$ is EQ1. ∎
See 4.4
@@
We give a brief algorithm description which is along the lines of Algorithm 1, except for the transfers which are more involved in this setting. We iteratively pick a poor agent $p$ and allocate it an item valued at $1$ or 00. If there is no such item for $p$, then we iteratively pick one of the rich agents and allocate it an item valued at $-1$ or 00. At this point, it is easy to see that the partial allocation $A$ is EQ1. As previously, we denote the set of agents with minimum utility under $A$ as poor agents $(P)$ and those with maximum utility under $A$ as rich agents ($R$). Either we arrive at a complete EQ1 allocation by continuing in the above manner, or there are unassigned items which are chores for agents in $P$ and goods for agents in $R$. Allocating any of the remaining items only increases the inequity and hence violates EQ1. Therefore, in order to move towards a complete EQ1 allocation, we now aim to convert a poor (rich) agent to rich (poor) by re-allocating one of the allocated items, so that the converted agent can now be a potential owner of one of the remaining unallocated items while maintaining EQ1. We do so by executing one of the following transfers at a time, after which at least one of them becomes a potential owner of $o\in U$.
@@
    Every $p$ values every good in $A_{r}$ at $-1$.
@@
    Every $r$ values every chore in $A_{p}$ at $1$
@@
    Every $p$ values every chore in $A_{p^{\prime}}$ at $-1$.
@@
Example A.2.
Leximin++ is not EQ1 even with $\{1,-1\}$-normalized valuations. The example in Table 6 illustrates this. The highlighted allocation is leximin++ with the utility vector as $\{3,1,1,1,1\}$. This is not EQ1 since even if the last $5$ agents choose to hypothetically ignore a good from $a_{1}^{\prime}s$ bundle, they still fall short of the equitable utility.
$o_{1}$ $o_{2}$ $o_{3}$ $o_{4}$ $o_{5}$ $o_{6}$ $o_{7}$ $o_{8}$ $o_{9}$ $o_{10}$ $o_{12}$ $o_{13}$
$a_{1}$1 1 1 1 1 1 -1 -1 -1 -1 -1 -1
$a_{2}$ 1 -1 -1 -1 -1 -1 -1 -1 1 1 1 1
$a_{3}$ 1 -1 -1 -1 -1 -1 -1 -1 1 1 1 1
$a_{4}$ 1 -1 -1 -1 -1 -1 -1 -1 1 1 1 1
$a_{5}$ 1 -1 -1 -1 -1 -1 -1 -1 1 1 1 1
$a_{6}$ 1 -1 -1 -1 -1 -1 -1 -1 1 1 1 1
Appendix B Omitted Details of Section 5
See 5.2
Proof.
We present a reduction from $3$-Partition, known to be strongly NP-hard, where the problem is to decide if a multiset of integers can be partitioned into triplets such that all of them add up to a constant. Formally, the input is a set $S=\{b_{1},b_{2},\ldots b_{3r}\}$; $r\in\mathbb{N}$; and the output is a partition of $S$ into $r$ subsets $\{S_{1},S_{2},\ldots S_{r}\}$ such that $\sum_{b_{i}\in S_{i}}b_{i}=T=\frac{1}{r}\sum_{b_{i}\in S}b_{i}$. Given any instance of $3$-Partition, we construct an instance of the allocation problem as follows. We create $r$ set-agents namely $\{a_{1},a_{2},\ldots a_{r}\}$ and one dummy agent, $a_{r+1}$. We create $3r$ set-items namely $\{o_{1},o_{2},\ldots o_{3r}\}$ and two dummy items $\{o_{3r+1},o_{3r+2}\}$. All the set agents value the set items identically at $b_{i}$ and the two dummy items at $-T$ and $-(r-3)T$ respectively. The dummy agent values the set-items $\{o_{1},\ldots o_{3r-1}\}$ at 00, $o_{3r}$ at $-(r-2)T$ and the two dummy items at $T$ and $(r-1)T$ respectively. Note that this is a type-normalized instance such that every agent values all the items together at $2T$. This completes the construction (see Table 4).
Forward Direction.
Suppose $3$-Partition is a yes-instance and $\{S_{1},S_{2},\ldots S_{r}\}$ is the desired solution. Then, the corresponding EQ1$+$PO allocation, $A$, can be constructed as follows. For $a_{i}:i\in[r]$, we set $A(a_{i})=\{o_{j}:o_{j}\in S_{i}\}$. Finally, $A(a_{r+1})=\{o_{3r+1},o_{3r+2}\}$. It is easy to see that $A$ is an EQ1 allocation since $v_{i}(A_{i}))=T$ for all $a_{i}:i\in[r]$ and $v_{i}(A_{i})\geq v_{r+1}(A(a_{r+1})\setminus\{o_{3r+2}\})=T$. Also, $A$ is PO since it is a welfare-maximizing allocation where each item is assigned to an agent who values it the most.
Reverse Direction.
Suppose that $A$ is an EQ1$+$PO allocation. Then, both the dummy items $\{o_{3r+1},o_{3r+2}\}$ must be allocated to the dummy agent $a_{r+1}$, who is the only agent who values both the items positively. (Since any set-agent derives a negative utility from these items, allocating them to its violates PO). Since $A$ is also EQ1, this forces every set-agent to derive the utility of at least $T$ under $A$ so that when it removes the item $\{o_{3r+2}\}$ from $A(a_{r+1})$, EQ1 is preserved. ∎
Algorithm 3 EQ1$+$PO, n=2 and $\{-1,0,1\}$ Type-Normalized Valuations
Input: An instance with $2$ agents, $m$ items and $\{1,0,-1\}$-type-normalized valuation function.
Output: An EQ1$+$PO allocation $A$.
1:  $A\leftarrow$ An empty allocation
@@ Appendix B Omitted Details of Section 5
(continued) We show here that if $A$ is not EQ1, then there is no EQ1$+$PO allocation for the instance. If the EQ1 violators in $A^{\prime}$ (the Nash allocation on the reduced instance) cannot be resolved by using the remaining chores, then there is no EQ1$+$PO allocation. We argue this claim by contradiction. Suppose there is a complete EQ1$+$PO allocation, say $A^{\star}$. Let $t$ be the number of items in $O^{-}$ that are valued at $-1$ by all the agents (they are ‘universal’ chores). Under the Nash optimal partial allocation $A^{\prime}$, let $v_{p}=\min_{i}v_{i}(A^{\prime}_{i})$ be the utility of the poorest agent, $v_{r}=v_{p}+1$ be the utility of rich agents, and the remaining agents are EQ1 violators with utility strictly greater than $v_{r}$. We denote the set of violators as $S$. It is easy to see that if $\sum_{s\in S}(v_{s}-v_{r})\leq t$, then the $t$ many $-1$’s from $O^{-}$ could have been used to bring down the utility of all the violators to $v_{r}$, ensuring that the completion of $A^{\prime}$ into $A$ is EQ1. Therefore, we have $\sum_{s\in S}(v_{s}-v_{r})>t$.
@@
Now consider the EQ1$+$PO allocation $A^{\star}$. Let $\bar{A^{\star}}$ denote the restriction of $A^{\star}$ to $O^{\pm}\cup O^{+}$. Then, $\bar{A^{\star}}$ must be a PO allocation but is not EQ1, otherwise the Nash optimal allocation must have been EQ1. Let $v_{p}^{\star},\leavevmode\nobreak\ v_{r}^{\star},\leavevmode\nobreak\ v_{s}^{\star}$ denote the utilities of the poor, rich, and violators in $\bar{A^{\star}}$, and $S^{\star}$ be the set of violators. It must be the case that $\sum_{v\in V^{\star}}(v_{v}^{\star}-v_{r}^{\star})\leq t$, since the completion $A^{\star}$ is an EQ1 allocation. Also, notice that since both $\bar{A^{\star}}$ and $A^{\prime}$ are Pareto optimal, they both allocate the items in $O^{\pm}\cup O^{+}$ non-wastefully. Therefore, the sum of the utilities under both the partial allocations is exactly $m^{\prime}$, where $m^{\prime}:=|O^{\pm}\cup O^{+}|$. We now consider the following cases.
  • $v_{p}\geq v_{p}^{\star}$. Then,
    $\sum_{p,\leavevmode\nobreak\ r,\leavevmode\nobreak\ s}(v_{p}+v_{r}+v_{s})>\sum_{p,\leavevmode\nobreak\ r}(v_{p}+v_{r})+(t+v_{r}|S|)$
    $\geq\sum_{p,\leavevmode\nobreak\ r}(v_{p}^{\star}+v_{r}^{\star})+(t+v_{r}^{\star}|S|)$
    $\geq\sum_{p,\leavevmode\nobreak\ r,\leavevmode\nobreak\ v}(v_{p}^{\star}+v_{r}^{\star}+v_{v}^{\star})$
    which is a contradiction as the first and the last term is equal to $m^{\prime}$.
  • $v_{p}<v_{p}^{\star}$. Then we will show that $A^{\prime}$ is not a Nash optimal allocation. We have $v_{r}=v_{p}+1<v_{p}^{\star}+1=v_{r}^{\star}$. Consider the nash welfare
    $NW(A^{\prime})=(\prod_{p,r,s}\leavevmode\nobreak\ v_{p}v_{r}v_{s})^{\frac{1}{n}}<(\prod_{v,r,s}\leavevmode\nobreak\ v_{p}^{\star}v_{r}^{\star}v_{s})^{\frac{1}{n}}$
    $<(\prod_{p,r,s}\leavevmode\nobreak\ v_{p}^{\star}v_{r}^{\star}v_{s}^{\star})^{\frac{1}{n}}=NW(\bar{A^{\star}})$
@@
    The last but one inequality follows from the fact that $\sum_{p,\leavevmode\nobreak\ r,\leavevmode\nobreak\ s}(v_{p}+v_{r}+v_{s})=\sum_{p,\leavevmode\nobreak\ r,\leavevmode\nobreak\ s}(v_{p}^{\star}+v_{r}^{\star}+v_{s}^{\star})=m^{\prime}$ and the product of a set of numbers with a constant sum has the highest outcome if they are closer/equal to each other rather than being further away ($\sum_{s\in V^{\star}}(v_{s}^{\star}-v_{r}^{\star})\leq t<\sum_{v\in V}(v_{s}-v_{r})$ ensures that $v_{r}^{\star}$ and $v_{s}^{\star}$ are closer to each other than $v_{s}$ and $v_{r}$).
This settles the claim. ∎
See 5.6

### v3

@@ Appendix B Omitted Details of Section 3
The algorithm iteratively picks a poor agent—breaking ties arbitrarily—and allocates it the most valuable good among the remaining items in $O^{+}$ according to its preference. Once $O^{+}$ is exhausted, it picks a rich agent who then receives its most disliked item from $O^{-}$.
@@
We show that Algorithm 1 returns an EQ1 allocation when the instance has $\{-1,1\}$ normalized valuations. First, note that since the instance is normalized, the reduced instance restricted to items in only $O^{\pm}$ is also normalized. That is, every agent assigns a value of $1$ to exactly $k_{1}$ items from $O^{\pm}$ and a value of 1-1 to exactly $k_{2}$ items from $O^{\pm}$, where $k_{1}$ and $k_{2}$ are constants. The idea is to first allocate all the items in $O^{\pm}$ in an almost equitable way and then to extend the partial EQ1 allocation to a complete allocation using Lemma 3.4. To that end, we first show that there exists a partial EQ1 allocation that allocates all items in $O^{\pm}$.
Lemma B.1.
@@
We first describe the algorithm. We iteratively pick a poor agent and allocate it an item it values at $1$ (a good) from $O^{\pm}$. If there are no such items in $O^{\pm}$, then we iteratively pick one of the rich agents and allocate it an item it values at 1-1 (a chore) from $O^{\pm}$. At this point, it is easy to see that the partial allocation is EQ1. We denote the set of agents with minimum utility as poor agents $(P)$ and those with maximum utility as rich agents ($R$). Now suppose that the remaining items in $O^{\pm}$ are all chores for the poor agents and goods for the rich agents, and if they are allocated to any of them, they increase the amount of inequity and hence may violate EQ1. Therefore, in order to move towards an EQ1 allocation that allocates all the items in $O^{\pm}$, we now aim to convert a poor (rich) agent to rich (poor) by re-allocating one of the allocated items, so that the converted agent can now be a potential owner of one of the remaining unallocated items while maintaining EQ1. We do so by executing one of the following transfers at a time, each of which allows us to allocate at least one unallocated item while respecting EQ1.
@@
    Every $p$ values every good in $A_{r}$ at 1-1.
@@
    Every $r$ values every chore in $A_{p}$ at 11
@@
    Every $p$ values every chore in $A_{p^{\prime}}$ at 1-1.
With these valuations at hand, we are now ready to prove Lemma B.1.
Proof of Lemma B.1.
@@
    If $|P|>|R|$, then we take a good ($1$-valued item) from $A_{r}$ and allocate it to $p$ who necessarily values it at 1-1 (else, there is an R-P transfer). We do this for $|R|$ disjoint pairs of a rich and a poor agent, thereby decreasing the utility of all these pairs by 1-1. But since $|P|>|R|$, we have a poor agent $p^{\prime}$ whose value remains intact, which in turn, makes him one of the rich agents after the above transfers. Now, the remaining unallocated item, which was a chore for $p^{\prime}$ can be allocated to it without violating EQ1.
@@
    Even if $p$ values every good in every other poor agent’s bundle at $1$, then also only $|P|k$ out of $|R|(k+1)+1$ such items are accounted for. (Note that $p$ can not value a chore in $A_{p^{\prime}}$ at $1$, else there exists a P-P transfer). Also, note that $p$ cannot value any of the $|R|(k+1)$ items allocated to the rich agents at $1$, else there is an R-P transfer. This implies that there must be at least (|R||P|)k+|R|+1(|R|-|P|)k+|R|+1 many allocated items valued at $1$ by $p$, outside of the $|R|(k+1)$ and $|P|k$ items allocated as goods to the rich and poor agents respectively. At this point, we say that $r$ stays ahead of $p$ in terms of the number of items valued as a good.
    Since $v_{r}(A_{r})=k+1$, in particular, $r$ derives its value of $k+1$ from a set of $k+1$ items valued at $1$ by itself. We call these $k+1$ items as contributing goods in the bundle $A_{r}$. There can be additional items in $A_{r}$, however, they do not contribute to the utility of $r$. Any such additional good $g$ (chore $c$) in $A_{r}$ is non-contributing, that is, it can be paired up with a chore $c$ (good $g$) in $A_{r}$ so that the final utility contribution of $g\cup c$ to $r$ is 0. We refer to these pairs as non-contributing good-chore pair $(g,c)\in A_{r}$. Likewise, for any poor agent $p$, there are $k$ contributing goods in $A_{p}$, and the rest of the goods (chores) in $A_{p}$ can be paired up with a chore (good) in $A_{p}$ so their cumulative utility contribution is 0.
@@
    For $p$ to compensate for this deficit, it must value at least (|R||P|)k+|R|+1(|R|-|P|)k+|R|+1 many allocated items, outside of the non-contributing pairs in any rich bundle, at $1$. Let’s call this set of goods for $p$ as $E_{g}^{p}$. Since we have counted for the contributing goods in any poor bundle (at most $|P|k$ of them), for any $g\in E_{g}^{p}$, we must have that it is allocated to a poor agent $p^{\prime}$ as a non-contributing pair. (Again, the non-contributing pairs allocated to $r$ are already counted for). Let $c$ be the corresponding non-contributing chore in $A_{p^{\prime}}$. If $v_{p}(c)=1$, then we have a P-P transfer. So, for all $p\in P$, we have $v_{p}(c)=-1$. Now, if any rich agent values $c$ at $1$, then it again stays ahead of the poor agents by this additional $1$ valued item (which is outside its bundle), and the process of compensating for the deficit does not terminate. Otherwise, $v_{r}(c)=-1$, but this is a contradiction to the fact that $c\in O^{\pm}$.
@@
$r_{1}$ 11 11 11 11 11 11 $\boxed{1,-1}$ 1-1,1-1 11 11
$r_{2}$ 11 11 11 11 11 11 1-1,1-1 11 11
$p_{1}$ 1-1 1-1 1-1 1-1 1-1 1-1 11 11 11 11 $\star 11,$ 1-1 1-1 1-1
$p_{2}$ 1-1 1-1 1-1 1-1 1-1 1-1 11 11 11 11 $\star 1,1$ 1-1 1-1
Table 6: A schematic of Case (2b) in the proof of Lemma B.1. $r_{1}$ and $r_{2}$ are two rich agents with utility 33 each and $p_{1}$ and $p_{2}$ are two poor agents with utility 22 each. The values in gray are forced, otherwise, there is a feasible transfer. The values in red depict the contradiction that there is an item $c$ such that $c\notin O^{\pm}$
@@
Lemma B.1 gives us a partial EQ1 allocation $A$. Now we can allocate the items from $O^{+}$ to a poor agent iteratively and once $O^{+}$ is exhausted, we allocate the item from $O^{-}$ to a rich agent iteratively (Lemma 3.4). This ensures that $A$ is EQ1 at every step of the algorithm. Consequently, the complete allocation $A$ is EQ1. ∎
See 4.3
@@
We give a brief algorithm description which is along the lines of Algorithm 1, except for the transfers, which are more involved in this setting. We iteratively pick a poor agent $p$ and allocate it an item valued at $1$ or 0. If there is no such item for $p$, then we iteratively pick one of the rich agents and allocate it an item valued at 1-1 or 0. At this point, it is easy to see that the partial allocation $A$ is EQ1. As previously, we denote the set of agents with minimum utility under $A$ as poor agents $(P)$ and those with maximum utility under $A$ as rich agents ($R$). Either we arrive at a complete EQ1 allocation by continuing in the above manner, or there are unassigned items which are chores for agents in $P$ and goods for agents in $R$. Allocating any of the remaining items only increases the inequity and hence violates EQ1. Therefore, in order to move towards a complete EQ1 allocation, we now aim to convert a poor (rich) agent to rich (poor) by re-allocating one of the allocated items, so that the converted agent can now be a potential owner of one of the remaining unallocated items while maintaining EQ1. We do so by executing one of the following transfers at a time, after which at least one of them becomes a potential owner of $o\in U$.
@@
    Every $p$ values every good in $A_{r}$ at 1-1.
@@
    Every $r$ values every chore in $A_{p}$ at 11
@@
    Every $p$ values every chore in $A_{p^{\prime}}$ at 1-1.
@@
Appendix C Omitted Details of Section 5
Algorithm 3 EQ1++PO, n=2 and $\{-1,0,1\}$ Type-Normalized Valuations
Input: An instance with $2$ agents, $m$ items and $\{1,0,-1\}$ type-normalized valuations.
Output: An EQ1++PO allocation $A$.
[…]
@@ Appendix C Omitted Details of Section 5
(continued) We show here that if $A$ is not EQ1, then there is no EQ1++PO allocation for the instance. If the EQ1 violators in $A^{\prime}$ (the Nash allocation on the reduced instance $\mathcal{I_{G}}$) cannot be resolved by using the remaining chores, then there is no EQ1++PO allocation. We argue this claim by contradiction. Suppose there is a complete EQ1++PO allocation, say $A^{\star}$. Let $t$ be the number of items in $O^{-}$, and hence valued at 1-1 by all the agents (they are ‘universal’ chores). Under the Nash optimal partial allocation $A^{\prime}$, let $v_{p}=\min_{i}v_{i}(A^{\prime}_{i})$ be the utility of the poorest agent and denote the set of all such agents as $P$, let the agents with utility $v_{r}=v_{p}+1$ be called the rich agents, denoted as $R$, and the remaining agents with utility strictly greater than $v_{r}$ be called the EQ1 violators, denoted by $S$. It is easy to see that if $\sum_{s\in S}(v_{s}-v_{r})\leq t$, then the $t$ many 1-1 valued chores from $O^{-}$ could have been used to bring down the utility of all the violators to $v_{r}$, ensuring that the completion of $A^{\prime}$ into $A$ is indeed EQ1. Therefore, we assume
$\sum_{s\in S}(v_{s}-v_{r})>t$ (4)
@@
Now consider the EQ1++PO allocation $A^{\star}$. Let $\overline{A^{\star}}$ denote the restriction of $A^{\star}$ to $O^{\pm}\cup O^{+}$. Then, $\overline{A^{\star}}$ must be a PO allocation but not EQ1, otherwise the Nash optimal allocation must have been EQ1 (Lemma 5.4). Let $v_{p}^{\star},\leavevmode\nobreak\ v_{r}^{\star},\leavevmode\nobreak\ v_{s}^{\star}$ denote the utilities of the poor agents $P$, rich agents $R$, and violators $S^{\star}$ in $\overline{A^{\star}}$. It must be the case that
$\sum_{s\in S^{\star}}(v_{s}^{\star}-v_{r}^{\star})\leq t$ (5)
since the completion $A^{\star}$ is an EQ1 allocation. Also, notice that since both $\overline{A^{\star}}$ and $A^{\prime}$ are Pareto optimal, they both allocate the items in $O^{\pm}\cup O^{+}$ non-wastefully. Therefore, the sum of the utilities under both the partial allocations is exactly $m^{\prime}$, where $m^{\prime}:=|O^{\pm}\cup O^{+}|$. Therefore,
$\sum_{p\in P,\leavevmode\nobreak\ r\in R,\leavevmode\nobreak\ s\in S}(v_{p}+v_{r}+v_{s})=\sum_{p\in P^{*},\leavevmode\nobreak\ r\in R^{*},\leavevmode\nobreak\ s\in S^{*}}(v_{p}^{\star}+v_{r}^{\star}+v_{s}^{\star})=m^{\prime}$ (6)
@@
Since all the $v_{i}$’s and $v_{i}^{\star}$’s are non-negative (because of non-wasteful allocation of items in $O^{\pm}\cup O^{+}$ under both $A^{\prime}$ and $\overline{A^{\star}}$), by AM-GM inequality [^5] For any $n$ non-negative numbers $\{v_{1},\ldots v_{n}\}$ arithmetic Mean is greater than the geometric mean (that is, $\frac{1}{n}\sum_{i\in[n]}{v_{i}}\geq(\prod_{i\in[n]}v_{i})^{\frac{1}{n}}$), it follows that the product of a set of numbers with a constant sum has the higher outcome if they are equal/closer to each other rather than being further away ($\sum_{s\in S^{\star}}(v_{s}^{\star}-v_{r}^{\star})\leq t<\sum_{s\in S}(v_{s}-v_{r})$ ensures that $v_{r}^{\star}$ and $v_{s}^{\star}$ are closer to each other than $v_{s}$ and $v_{r}$ (Equations 4 and 5)). Therefore,
$NW(A^{\prime})=\left(\prod_{p\in P,r\in R,s\in S}\leavevmode\nobreak\ v_{p}v_{r}v_{s}\right)^{\frac{1}{n}}<\left(\prod_{p\in P^{*},r\in R^{*},s\in S^{*}}\leavevmode\nobreak\ v_{p}^{\star}v_{r}^{\star}v_{s}^{\star}\right)^{\frac{1}{n}}=NW(\overline{A^{\star}})$ (7)
Hence, $A^{\prime}$ is not a Nash optimal allocation. This settles the claim. ∎
See 5.7